the point of intersection of the pair of the pair of lines represented by xy+x+y+1=0 is

Grade:9

1 Answers

Vijay Mukati
8 years ago
xy+x+y+1=0 can be written as (x+1)(y+1)=0. Which represent two lines X+1=0 and Y+1=0.And therefore the point of intersection is (-1,-1).
Thanks.